OpenStudy (anonymous):

What are the amplitude and period of the function: f(x) = 4 + 8cos(7x)cos(10) - 11sin(9x)sin(8)?

OpenStudy (anonymous):

4+8cos7xcos10x-11sin9xsin8x

OpenStudy (anonymous):

4+4cos17x+4cos3x+5.5cos17x-5.5cosx=4+9.5cos17x+4cox3x-5.5cosx T=[2pi/17,2pi/3,2pi]=2pi/51[3,17,51]=2pi*51/51=2pi you cannot determine the amplitude so it is more than 4+9.5+4-5.5=12 and less than 4+9.5+4+5.5=23

OpenStudy (anonymous):

if you want to get exact amplitude you must take derivation so it is going to get hard!!

OpenStudy (anonymous):

Cool, thanks so much for your help.